Circular No. 8012

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002hp, 2002hq, 2002hr, 2002hs, 2002ht, 2002hu, 2002hv

Further to IAUC 7981, A. Riess, Space Telescope Science Institute, on behalf of GOODS Team (+ N. Grogin, A. Hornschemeier, R. Lucas, and M. Richardson) and the Hubble Higher-z Supernova Team, reports the discovery of five additional apparent supernovae on CCD images taken during Oct. 31–Nov. 3 with the Hubble Space Telescope (+ ACS) during the GOODS observations of the Chandra Deep Field South. Each new object was detected independently in each of four separate z-band (F850lp) exposures, via subtraction of images taken during Sept. 19–22. None of objects was apparent in earlier ACS z-band imaging of the fields taken July 31–Aug. 4. Four of the five objects were detected also in two i-band (F775W) exposures, and three in two v-band (F606W) exposures. The offsets given below are to the center of the nearest visible galaxy.

SN  2002 UT  α_{2000}  δ_{2000}  z  Offset
2002hp Nov. 1.51 3^{h}32^{m}24.^{s}79 −27^{o}46′17′′.8 24.4 0′′.01 W, 0′′.02 N
2002hq  1.51 3 32 30.12 −27 43 46.7 25.7 0′′.91 W, 0′′.18 S
2002hr  1.64 3 32 22.57 −27 41 52.2 23.5 0′′.03 E, 0′′.05 N
2002hs  2.23 3 32 18.59 −27 48 33.7 25.6 0′′.70 W, 0′′.35 S
2002ht  2.51 3 32 09.32 −27 41 29.3 25.5 0′′.34 E, 0′′.48 N

Reference z-band limiting magnitudes: SN 2002hp, Sept. 20.33 UT, [27.5; SN 2002hq, Sept. 20.19, [27.5; SN 2002hr, Sept. 20.99, [28; SN 2002hs, Sept. 20.51, [28; SN 2002ht, Sept. 20.51, [28.

Two more apparent supernovae have been reported on unfiltered CCD images: SN 2002hu in MCG +06-6-12 by T. Boles (cf. IAUC 8011); and 2002hv in UGC 4974 by LOTOSS with the 0.8-m Tenagra II automated telescope, via M. Schwartz and W. Li (cf. IAUC 7906).

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002hu Nov. 7.09 2^{h}18^{m}20.^{s}06 +37^{o}27′58′′.8 17.0 20′′.6 E, 6′′.6 N
2002hv  5.5 9 22 10.91 +33 50 10.8 18.6 7′′.1 E, 43′′.2 S

Boles obtained a confirming image of SN 2002hu on Nov. 7.815 UT; it was not present on Boles’ images from Oct. 4 (limiting mag 18.5) or Sept. 8, and it is not present on Palomar Sky Survey red and blue plates. Schwartz obtained a confirming image of SN 2002hv on Nov. 7.5, while an unfiltered KAIT CCD image taken at Lick Observatory on Apr. 28.2 showed nothing at that position (limiting mag ∼ 19.0).

2002 November 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green